OpenStudy (anonymous):

Five added to twice Erik's age is the same as 3 times his age minus 2. How old is Erik?

OpenStudy (anonymous):

5+2x=3x-2 solve for x

OpenStudy (anonymous):

5+2x=3x-2 subtract 2x from each side 5=x-2 add 2 to both sides x=7

OpenStudy (anonymous):

i knew it

OpenStudy (anonymous):

Ha ha well good job then